Will rate!!





Assuming that interest rate is 5%, solve for w in the CFD below. 150 300 45o 600 750 150 450A Enter your answer as follow: 123456.78. Do not use any symbol.
Answer #1


Calculate the value of W -

W = 150(F/P, 5%, 2) + 300(F/P, 5%, 1) + 450 + 600(P/F, 5%, 1) + 750(P/F, 5%, 2)

W = (150 * 1.1025) + (300 * 1.0500) + 450 + (600 * 0.9524) + (750 * 0.9070)

W = 165.37 + 315 + 450 + 571.44 + 680.25

W = 2182.06

The value of W is 2182.06